OpenStudy (anonymous):

10+x/4= x-11/3(show work)? 10+x/4= x-11/3(show work)? @Mathematics

OpenStudy (anonymous):

multiply by 12: 120+3x=4(x-11) 120+3x=4x-44 120+44=x x=164

OpenStudy (anonymous):

10 + x/4 = x - 11/3 10 + 11/3 = x - x/4 41/3 = (3x)/4 4x41/3 = 3x (4x41)/(3x3) =x x = 164/9

OpenStudy (anonymous):

wee need to know exactly where the bracket ( , ) is, otherwise we may come up with many solution :)

OpenStudy (anonymous):

true :P

OpenStudy (anonymous):

none of these answers match my choices. The problem is finding x by dvision

OpenStudy (anonymous):

is it x-11 or x - 11 ? ---- --- 3 3